OpenStudy (anonymous):

Determine whether y varies directly with x. if so, find the constant of variation k and write the equation. Ill include the table, but I just need to know what formula to use.

OpenStudy (anonymous):

I know the formula is either y=kx or k=y/x. I tried k=y/x but I didn't get the same answers for each pair.

OpenStudy (anonymous):



jimthompson5910 (jim_thompson5910):

you use y = k*x

jimthompson5910 (jim_thompson5910):

you can solve for k to get k = y/x if you want

jimthompson5910 (jim_thompson5910):

notice how for the first row we get k = y/x k = 11/7 and for the second row we get k = y/x k = 13/8 but those two values aren't the same, so this means that y does NOT vary directly with x

OpenStudy (anonymous):

I haven't used that formula before. For example, for the first pair, do I: 11=k*7 ?

jimthompson5910 (jim_thompson5910):

then you solve for k

OpenStudy (anonymous):

So, since y doesn't vary directly with x I cant continue doing the problem?

jimthompson5910 (jim_thompson5910):

that is correct k is not the same each time, so y doesn't vary directly with x and because of that you can't find the equation in the form y = k*x

OpenStudy (anonymous):

I see. so, to show my work Im going to solve each pair using the formula k=y/x and after Im going to say y does not vary directly with x so I cant find the equation in the form of y=k*x. ?

jimthompson5910 (jim_thompson5910):

yes all you need are 2 rows or two pairs to show that k is different

jimthompson5910 (jim_thompson5910):

you don't need to work with every pair unless you want to

OpenStudy (anonymous):

Alright, thank you!

jimthompson5910 (jim_thompson5910):

you're welcome